If cube root of (12x + 28)= 4 what is the value of x^3 ?
(A) -8 (B) 3 (C) 12 (D) 27 (E) 64

Answer by mananth(16946)

cube root of (12x + 28)= 4
Take cube of both sides
12x+28= 64
12x=64-28
12x=36
x= 36/6
x=3